What is the pH of a buffer solution made by mixing 0.1 M benzoic acid and 0.15 M of its conjugate base, sodium benzoate. Ka = 6.5x10-5 for benzoic acid.
Answer #1

We know pKa = - log Ka = - log (6.5 * 10^-5) = - log 6.5 + 5 log 10 = - 0.81 + 5 * 1 = 4.19

We know from Henderson equation

pH = pKa + log {[conjugate base] / [Acid]}

pH = 4.19 + log ( 0.15 M / 0.1 M)

pH = 4.19 + log 1.5 = 4.19 + 0.176 = 4.37

pH = 4.37